# No Imputation of Missing Values In Tabular Data Classification Using Incremental Learning

### v1

@@ Abstract
Tabular data sets with varying missing values are prepared for machine learning using an arbitrary imputation strategy. Synthetic values generated by imputation models often concern data stakeholders about computational complexity, data quality, and data-driven outcomes. This paper eliminates these concerns by proposing no imputation incremental learning (NIIL) of tabular data with varying missing value rates and types. The proposed method incrementally learns partitions of overlapping feature sets while using attention masks to exclude missing values from attention scoring. The average classification performance rank order across 15 diverse tabular data sets highlights the superiority of NIIL over 11 state-of-the-art learning methods with or without missing value imputations. Further experiments substantiate the robustness of NIIL against varying missing value types and rates compared to methods that involve the imputation of missing values. Our empirical analysis reveals that a feature partition size of half of the original feature space is, computation-wise and accuracy-wise, the best choice for the proposed incremental learning. The proposed method is one of the first deep learning solutions that can effectively learn tabular data without requiring the imputation of missing values.
@@ 2.2 Proposed NIIL algorithm
Head ={($\text{Head =}\begin{cases}(0*I_{d\times h})*V,&\text{if $M_{1}(:)$\text{ = -$\infty$}, $M_{2}(:)=0$}\\ \text{softmax}(\frac{QK^{T}}{\sqrt{h}})V,&\text{if $M_{1}(:)=0$, $M_{2}(:)=1$}\end{cases}$ (6)
$M_{1}$ uses $\infty$- ∞ to mask the attention columns with missing features. The remaining rows of attention with missing features are masked by 0s of $M_{2}$. Otherwise, $M_{1}$ and $M_{2}$ retain the attention scores of all observed feature pairs using 0s and 1s, respectively.
@@ 3.1 Tabular data sets
Tabular data sets with large sample sizes ($>$10,000) are often selected to facilitate deep learning [35, 36]. However, most tabular data sets in practice have relatively smaller sample sizes ($<$1000) [37]. For generalizability, we have selected 15 diverse data sets with a variety of numerical and categorical feature mixtures from the OpenML [38] repository. The sample size of these data sets ranges from 155 to 2109, while the feature dimension is between 7 and 39. A summary of the tabular data sets used in this article is in Table 1, where the first 13 data sets are available in complete form without missing values. The remaining two data sets (IDs 55, and 6332) have missing values that occur naturally at the rate of 48.8% and 48.7%, respectively.
@@ 4.4 Ablation studies
In an ablation study, we vary the value of feature partition size k for incremental learning using three representative data sets: Kc2 (ID 1063), Diabetes (ID 37), and Dresses sales (ID 23381). Figure 4 shows the k values that yield the best AUC scores for individual data sets. In general, feature incremental learning is superior to learning with all features at a time (i.e., no incremental learning, k = d). The best average AUCs are obtained for the data sets Kc2 $(d=21)$ at k=10, Diabetes $(d=8)$ at k=6, and Dresses sales $(d=12)$ at k = 6. Therefore, the value of k can be chosen as half of the feature dimension for the best classification performance. This observation aligns with the computational cost analysis presented in Figure 5 in  A, where $k<\frac{d}{2}$ is computationally more efficient than using the entire feature space ($d$).NIIL outperforms all or most of the baselines at k = 3, which can be further improved by increasing the partition size to $\frac{d}{2}$. The proposed method is evaluated on tabular data sets with categorical-only, numerical-only, and mixed data types. The sample size of these data sets ranges from 155 to 2109. Most tabular data in practice have limited sample sizes ($<1000$) with heterogeneous data types. The imputation of data with limited samples poses an additional challenge. In contrast, deep learning studies selectively use tabular data sets with large sample sizes ($>10000$) only [35, 8]. Therefore, our results are presented using the most practical tabular data sets, missing types, and rates to ensure generalizability.

### v2

# Imputation Not Required in Incremental Learning of Tabular Data with Missing Values

@@ Abstract
Tabular data sets with varying missing values are prepared for machine learning using an arbitrary imputation strategy. Synthetic values generated by imputation models often raise concerns among data stakeholders about computational complexity, data quality, and data-driven outcomes. This paper addresses these concerns by proposing no-imputation incremental learning (NIIL) of tabular data with varying missing value rates and types. The proposed method incrementally learns partitions of overlapping feature sets while using attention masks to exclude missing values from attention scoring. The average classification performance rank order across 15 diverse tabular data sets highlights the superiority of NIIL over 11 state-of-the-art learning methods with or without missing value imputations. Further experiments substantiate the robustness of NIIL against varying missing value types and rates compared to methods that involve the imputation of missing values. Our empirical analysis reveals that a feature partition size of half the original feature space is, both computationally and in terms of accuracy, the best choice for the proposed incremental learning. The proposed method is one of the first deep learning solutions that can effectively learn tabular data without requiring the imputation of missing values.
@@ 1.2 Motivation and contributions
This article makes five key contributions to address the challenges in learning tabular data with missing values. First, we present one of the first studies to streamline tabular data with missing values in a deep attention learning framework without requiring imputation or even initialization of missing values. Second, the proposed method derives and retrofits two attention masks to a feature- tokenized transformer to exclude missing entries from attention scoring. Third, we propose a new feature incremental learning strategy to enable efficient attention learning of O ($n^{2}$), which otherwise results in an out-of-memory problem in large feature spaces [32]. Fourth, the proposed incremental learning utilizes data in small and fixed-size feature partitions, rather than gradually increasing feature sizes. A fixed and small partition allows for more efficient learning of a large-dimensional feature space than conventional feature incremental learning methods. Fifth, the entire deep attention framework is designed to be computationally efficient and outperform state-of-the-art baseline methods, both with and without missing value imputations.
@@ 2.2 Proposed NIIL algorithm
$\text{Head =}\begin{cases}(0*I_{d\times h})*V,&\text{if $M_{1}(:)$\text{ = -$\infty$}, $M_{2}(:)=0$}\\ \text{softmax}(\frac{QK^{T}}{\sqrt{h}})V,&\text{if $M_{1}(:)=0$, $M_{2}(:)=1$}\end{cases}$ (6)
$M_{1}$ uses -$\infty$ to mask the attention columns with missing features. The remaining rows of attention with missing features are masked by 0s of $M_{2}$. Otherwise, $M_{1}$ and $M_{2}$ retain the attention scores of all observed feature pairs using 0s and 1s, respectively.
@@ 3.1 Tabular data sets
Tabular data sets with large sample sizes (>>10,000) are often selected to facilitate deep learning [35, 36]. However, most tabular data sets in practice have relatively smaller sample sizes (<<1000) [37]. For generalizability, we have selected 15 diverse data sets with a variety of numerical and categorical feature mixtures from the OpenML [38] repository. The sample size of these data sets ranges from 155 to 2109, while the feature dimension is between 7 and 39. A summary of the tabular data sets used in this article is in Table 1, where the first 13 data sets are available in complete form without missing values. The remaining two data sets (IDs 55, and 6332) have missing values that occur naturally at the rate of 48.8% and 48.7%, respectively.
@@ 4.4 Ablation studies
In an ablation study, we vary the value of feature partition size k for incremental learning using three representative data sets: Kc2 (ID 1063), Diabetes (ID 37), and Dresses sales (ID 23381). Figure 4 shows the k values that yield the best AUC scores for individual data sets. In general, feature incremental learning is superior to learning with all features at a time (i.e., no incremental learning, k = d). The best average AUCs are obtained for the data sets Kc2 $(d=21)$ at k=10, Diabetes $(d=8)$ at k=6, and Dresses sales $(d=12)$ at k = 6. Therefore, the value of k can be chosen as half of the feature dimension for the best classification performance. This observation aligns with the computational cost analysis presented in Figure 5 in  A, where $k<\frac{d}{2}$ is computationally more efficient than using the entire feature space ($d$).NIIL outperforms all or most of the baselines at k = 3, which can be further improved by increasing the partition size to $\frac{d}{2}$. The proposed method is evaluated on tabular data sets with categorical-only, numerical-only, and mixed data types. The sample size of these data sets ranges from 155 to 2109. Most tabular data in practice have limited sample sizes (<1000<1000) with heterogeneous data types. The imputation of data with limited samples poses an additional challenge. In contrast, deep learning studies selectively use tabular data sets with large sample sizes (>10000>10000) only [35, 8]. Therefore, our results are presented using the most practical tabular data sets, missing types, and rates to ensure generalizability.